OpenStudy (anonymous):

Use substitution method to solve y = x^2 y = x + 2

hero (hero):

x^2 = x+2 x^2-x - 2 = 0 (x-2)(x+1) = 0 x = 2,-1

OpenStudy (anonymous):

so x=2, but y=-1 or 1?

OpenStudy (lilg132):

i think his answer is correct

hero (hero):

Yours was correct, you just solved for y. I solved for x

OpenStudy (lilg132):

well y = 1 then

hero (hero):

Two different approaches to the same goal

hero (hero):

Yes, y = 1

OpenStudy (anonymous):

oh ok

hero (hero):

But x can be -1 or 2

hero (hero):

y = 1 or 4